Given a market risk premium of 7.17% and a risk-free rate of 1.06%, what is the reward-to-risk ratio of a stock with a beta of 1.16 according to the CAPM? (answer as a fraction, give four decimal places)
Answer #1

Given,

Market risk premium = 7.17% or 0.0717

Risk free rate = 1.06% or 0.0106

Beta = 1.16

Solution :-

Expected return = risk free rate + (beta x market risk premium)

= 0.0106 + (1.16 x 0.0717)

= 0.0106 + 0.083172 = 0.093772

Reward-to-risk ratio = (expected return - risk free rate)/beta   

= (0.093772 - 0.0106)/1.16

= 0.083172/1.16 = 0.0717